# Read Between the Layers: Leveraging Intra-Layer Representations for Rehearsal-Free Continual Learning with Pre-Trained Models

### v2

@@ Abstract
We address the Continual Learning (CL) problem, wherein a model must learn a sequence of tasks from non-stationary distributions while preserving prior knowledge upon encountering new experiences. With the advancement of foundation models, CL research has pivoted from the initial learning-from-scratch paradigm towards utilizing generic features from large-scale pre-training. However, existing approaches to CL with pre-trained models primarily focus on separating class-specific features from the final representation layer and neglect the potential of intermediate representations to capture low- and mid-level features, which are more invariant to domain shifts. In this work, we propose LayUP, a new prototype-based approach to continual learning that leverages second-order feature statistics from multiple intermediate layers of a pre-trained network. Our method is conceptually simple, does not require access to prior data, and works out of the box with any foundation model. LayUP surpasses the state of the art in four of the seven class-incremental learning benchmarks, all three domain-incremental learning benchmarks and in six of the seven online continual learning benchmarks, while significantly reducing memory and computational requirements compared to existing baselines. Our results demonstrate that fully exhausting the representational capacities of pre-trained models in CL goes well beyond their final embeddings. The code will be made publicly available upon acceptance.
@@ 1 Introduction
Such methods for downstream adaptation can be categorized into two general strategies (Wang et al., 2023a; McDonnell et al., 2023): (i) carefully fine-tuning the pre-trained representations (full body adaptation), or (ii) keeping the pre-trained backbone fixed while learning a small set of additional parameters such as prompts (Lester et al., 2021; Liu et al., 2023; Jia et al., 2022), adapters (Rebuffi et al., 2017a; Houlsby et al., 2019; Hu et al., 2022; Chen et al., 2022; 2023), or other methods (Ke et al., 2021; Marullo et al., 2023; Lian et al., 2022). This strategy is commonly referred to as parameter-efficient transfer learning (PETL). Both strategies (i) and (ii) have in common that they do not rely on rehearsal from past data to perform well, unlike many of the top-performing CL strategies used for models trained from scratch. Despite their merits, most approaches to CL with pre-trained models that fall under one of those strategies update a fully connected classification head (linear probe) via iterative gradient methods during training and inference, which is subject to task-recency bias (Mai et al., 2021; Rymarczyk et al., 2023; Wang et al., 2023b) and catastrophic forgetting (Zhang et al., 2023; Ramasesh et al., 2021). Class-prototype methods (Zhou et al., 2023b; Janson et al., 2022; McDonnell et al., 2023) have been established as a viable alternative, as they utilize extracted features directly, which is not only computationally efficient but also superior to CL methods applied to the trainable linear probe.
@@
In this work, we hypothesize that representations are most effectively leveraged from multiple layers to construct a classifier based on first-order (class prototypes) and second-order (Gram matrix) feature statistics. Such a strategy is inspired by works on Neural Style Transfer (Gatys et al., 2016; Jing et al., 2020), which involve applying the artistic style of one image to the content of another image, a problem that can be viewed through the lens of domain adaptation (Li et al., 2017). To achieve neural style transfer, a model has to disentangle image information into content and style (e.g., textures, patterns, colors). While content information is expressed in terms of activations of features at different layers (which translates to calculating intra-layer class prototypes in our case), style information is expressed as correlations between features of different layers (which corresponds to calculating intra-layer Gram matrices).
@@
We thus propose Intra-Layer Universal Prototypes (LayUP), a class-prototype method for CL that is based on the aforementioned strategy to disentangle style and content information of images at multiple layers of a pre-trained network. An overview of our method is given in Fig. 3. LayUP obtains rich and fine-grained information about images to perform a more informed classification that is less sensitive to domain shifts. We additionally experiment with different parameter-efficient fine-tuning strategies to further refine the intra-layer representations obtained by LayUP. Across various CL benchmarks and learning settings, our method enhances performance and narrows the gap to the upper bound by as much as 80% (Stanford Cars-196 in the CIL setting) compared to the next best baseline, while also reducing its memory and compute requirements by 81% and up to 90%, respectively.
@@
Our contributions are threefold: (1) We present and examine in detail the CL strategy with pre-trained models and show why it benefits from using intra-layer representations for classification. We further demonstrate the advantages of leveraging cross-correlations between intra-layer and inter-layer features to decorrelate class prototypes. (2) Building on our insights, we propose LayUP, a novel class-prototype approach to CL that leverages second-order statistics of features from multiple layers of a pre-trained model. Inspired by prior works (Zhou et al., 2023a; McDonnell et al., 2023), we experiment with different methods for parameter-efficient model tuning and extend our method towards first session adaptation. Our final approach is not only conceptually simple, but also benefits from low memory and computational requirements, and integrates seamlessly with any pre-trained model. (3) We report performance improvements with a pre-trained Vit-B/16 (Dosovitskiy et al., 2021) backbone on the majority of benchmarks the Class-Incremental Learning (CIL) and Domain-Incremental Learning (DIL) settings (van de Ven & Tolias, 2019) as well as in the challenging Online Continual Learning (OCL) setting. We show that LayUP is especially effective under large distributional shifts and in the low-data regime. Our results highlight the importance of taking a deeper look at the intermediate layers of pre-trained models to better leverage their powerful representations, thus identifying promising directions for CL in the era of large foundation models.
@@ 2.1 Continual learning
Traditional approaches to CL consider training a model from scratch on a sequence of tasks, while striking a balance between adapting to the currently seen task and maintaining high performance on previous tasks. They can be broadly categorized into the paradigms of regularization (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2018; Aljundi et al., 2018; Dhar et al., 2019), which selectively restricts parameter updates, (pseudo-)rehearsal (Rebuffi et al., 2017b; Chaudhry et al., 2019; Buzzega et al., 2020; Prabhu et al., 2020), which recovers data either from a memory buffer or from a generative model, and dynamic architectures (Rusu et al., 2016; Yoon et al., 2018; Serra et al., 2018), which allocate disjoint parameter spaces for each task.
@@ 2.2 Continual learning with pre-trained models
As an alternative to the methods above that adjust the feature space via cross-entropy loss and backpropagation, constructing prototypes directly from the pre-trained features can avoid catastrophic forgetting during CL without storing data. Janson et al. (2022) and Zhou et al. (2023b) find that a conceptually simple Nearest Mean Classifier (NMC) outperforms various prompt learning methods. ADAM (Zhou et al., 2023b) combines NMC with FSA and concatenates the features from the initial pre-trained ViT and the adapted parameters. Other methods leverage second-order feature statistics, i.e., covariance information, for class prototype accumulation: FSA+LDA (Panos et al., 2023) applies an incremental version of linear discriminant analysis to a pre-trained ResNet encoder (He et al., 2016). RanPAC (McDonnell et al., 2023) uses high-dimensional random feature projections to decorrelate class prototypes of a pre-trained ViT backbone.
All the aforementioned approaches have in common that they consider only the last layer representations (or, features) for classification. We provide a different perspective to the prototype-based approach to CL and show that intra-layer representations can add valuable information to the linear transformation of class prototypes (cf. Eq. 5), which leads to better class separability and increased robustness to large distributional shifts.
@@ 3.1 Continual learning problem formulation
We consider a feature extractor $\phi(\cdot)$ composed of $L$ consecutive layers and a sequence of training datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$, where the $t$th task $\mathcal{D}_{t}=\{({\bm{x}}_{t,n},y_{t,n})\}_{n=1}^{N_{t}}$ contains pairs of input samples ${\bm{x}}_{t,n}\in\mathcal{X}_{t}$ and their ground-truth labels $y_{t,n}\in\mathcal{Y}_{t}$. The total number of classes seen in $\mathcal{D}$ (i.e., the number of unique elements in $\mathcal{Y}$) is denoted with $C$. We represent the $d_{l}$-dimensional encoded features from the $l$th layer of model $\phi$ from some test image ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$ with unknown label during inference as $\phi_{l}({\bm{x}})\in{\mathbb{R}}^{d_{l}}$. We consider rehearsal-free CL, where historical data cannot be fetched for replay. In the Class-Incremental Learning (CIL) setting, label spaces are partially disjoint between tasks. Conversely, in the Domain-Incremental Learning (DIL) setting, label spaces are kept across tasks, but samples within a label space are subject to distributional shifts. Online Continual Learning (OCL) is a more tightened definition of the aforementioned settings towards each datum in the stream of tasks being observed only once. All three CL settings prohibit the access of the model to task identifiers during testing (notably, our method does not require access to task-specific information during training either).
@@ 3.2 Class-prototype methods for CL with pre-trained models
To leverage the powerful representations from large-scale pre-training while overcoming the limitations of full body adaptation and linear probing, class-prototype methods construct representatives for each class from the last layer features of a pre-trained model. The most straightforward class-prototype method is the Nearest Mean Classifier (NMC), which aggregates per-class training sample features via averaging (denoted as $\overline{{\bm{c}}}_{y}$ for each class $y$) and, during inference, selects for each test sample’s feature vector the class with the smallest Euclidean distance (Janson et al., 2022) or highest cosine similarity (Zhou et al., 2023b) with its class prototype. Considering the cosine similarity measure and some ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, the predicted class label is obtained by:
@@ 4 LayUP
Figure 4: Classification performance of intermediate layers. Comparison for two different pre-training paradigms of a ViT-B/16 (Dosovitskiy et al., 2021) backbone and split ImageNet-R and CIFAR-100 datasets. For each intermediate layer $l\in\{1,\dots,L-1\}$ (where $L=12$ denotes the final representation layer), the bars represent the percentage of classes for which a classifier, utilizing Eq. 2 at layer $l$, surpasses the accuracy of the classifier at the $L$-th layer.
@@ 4.1 Class prototyping from second-order intermediate features
We argue that a combination of (i) enriching the last layer representations with hierarchical intra-layer features and (ii) decorrelating intra-layer features, which represent image properties such as content and style, via Gram matrix transformation increases robustness to domain shifts and thus improves generalizability to downstream continual tasks. To this end, we propose to integrate embeddings from multiple layers via concatenation into a ridge regression estimator as defined in Eq. 2. Let
@@
denote the concatenated input features of some input sample ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, extracted from the last $k$ layers of the pre-trained model $\phi(\cdot)$. Such features can be, e.g., class embeddings ([CLS]) of a transformer (Vaswani et al., 2017) encoder or flattened feature maps of a ResNet (He et al., 2016) encoder. This yields a modified estimator
@@
with $(d_{L-k+1}+\dots+d_{L-1}+d_{L})$-dimensional identity matrix ${\bm{I}}$ and
@@
We first motivate our approach by showing that intra-layer representations capture expressive class statistics for prototype-based classification. For the split CIFAR-100 (Krizhevsky, 2009) and ImageNet-R (Hendrycks et al., 2021a) datasets and two ViT models, we construct one prototype-based classifier using Eq. 2 per layer $l\in\{1,\dots,L-1\}$ and measure the percentage of classes that layer $l$ predicts better than the last layer $L$ ($L=12$ for ViT-B/16). The results are provided in Fig. 4. Across both datasets, the classifiers from the last five intermediate layers outperform the representation layer classifier in up to 32% of the classes. The advantage of intra-layer classifiers is more pronounced in the ViT model pre-trained on 1 000 different ImageNet classes (IN1K) compared to the ViT model pre-trained on 21 000 different ImageNet classes (IN21K). The results indicate that representations derived from intermediate layers generally provide meaningful knowledge to more effectively distinguish classes at various hierarchical levels.
@@
Figure 5: Comparison of techniques to integrate intra-layer representations. LayUP implementations for different values of $k$ using a shared representation and Gram matrix as in Eq. 5 versus averaging over separate ridge (or, Gram) classifiers for each layer using Eq. 2. Results are reported as average accuracy scores over CL training on split ImageNet-R (left) and CIFAR-100 (right) datasets following phase B of Alg. 1.
@@
There are two intuitive ways to integrate intra-layer representations into the class-prototype classifier: Averaging over $k$ separate classifiers per Eq. 2 or concatenating representations from the last $k$ layers to obtain shared class prototypes per Eq. 5. Fig. 5 shows average accuracies for two split datasets and different values of $k$. For each $k$, transforming shared class-specific information via Gram matrix inversion, as described in Eq. 5, proves to be superior to averaging across separate classifiers. This finding indicates that the correlations across layers, captured in the shared Gram matrix (cf. Eq. 6), add meaningful information to enhance class separability.
@@
While performance generally benefits from higher values of $k$, the difference in performance between $k=4$ and $k=6$ is more pronounced than that between $k=6$ and $k=12$. Considering that a higher $k$ corresponds to increased computational and memory demands as $\mathbf{G}$ and $\mathbf{c}_{y}$ increase in dimension, this substantiates $k=6$ as a reasonable choice for maximum representation depth. Although concatenating representations from multiple consecutive layers results in higher memory demands compared with using a last layer classifier or averaging over layer-wise classifiers, we will demonstrate in Sec. 4.3 that our method remains significantly lighter in memory and computational requirements than other competitive class-prototype methods for CL.
@@
Interestingly, the performance of the last-layer classifier $k=1$, as illustrated in Fig. 5, is approximately equivalent to that achieved by averaging over separate classifiers for $k=4$ and $k=6$. This observation challenges the argument that class embeddings from multiple consecutive layers, which are individually strong at predicting specific classes (cf. Fig. 4), might introduce noise that impairs performance when combined via averaging.

### v3

# Read Between the Layers: Leveraging Multi-Layer Representations for Rehearsal-Free Continual Learning with Pre-Trained Models

@@ Abstract
We address the Continual Learning (CL) problem, wherein a model must learn a sequence of tasks from non-stationary distributions while preserving prior knowledge upon encountering new experiences. With the advancement of foundation models, CL research has pivoted from the initial learning-from-scratch paradigm towards utilizing generic features from large-scale pre-training. However, existing approaches to CL with pre-trained models primarily focus on separating class-specific features from the final representation layer and neglect the potential of intermediate representations to capture low- and mid-level features, which are more invariant to domain shifts. In this work, we propose LayUP, a new prototype-based approach to CL that leverages second-order feature statistics from multiple intermediate layers of a pre-trained network. Our method is conceptually simple, does not require access to prior data, and works out of the box with any foundation model. LayUP surpasses the state of the art in four of the seven class-incremental learning benchmarks, all three domain-incremental learning benchmarks and in six of the seven online continual learning benchmarks, while significantly reducing memory and computational requirements compared to existing baselines. Our results demonstrate that fully exhausting the representational capacities of pre-trained models in CL goes well beyond their final embeddings.
@@ 1 Introduction
Recent approaches to CL with pre-trained models primarily adopt three main strategies (Wang et al., 2024; McDonnell et al., 2023): (i) carefully fine-tuning the parameters of the pre-trained model (full body adaptation), (ii) keeping the pre-trained parameters fixed while learning a small set of additional parameters such as prompts (Lester et al., 2021; Liu et al., 2023; Jia et al., 2022), adapters (Rebuffi et al., 2017a; Houlsby et al., 2019; Hu et al., 2022; Chen et al., 2022; 2023), or others (Ke et al., 2021; Marullo et al., 2023; Lian et al., 2022), or (iii) extracting class prototypes from pre-trained representations without further fine-tuning. All three strategies share the beneficial property of not relying on rehearsal from past data to perform well, unlike many of the top-performing CL strategies used for models trained from scratch. The optimal strategy for balancing the trade-offs between effectively utilizing pre-trained features, ensuring resource efficiency, and maintaining robustness to domain shifts in CL settings remains an open question.
Full body adaptation, as done with strategy (i), is expensive with respect to computational resources and required training data. Parameter-Efficient Transfer Learning (PETL), as implemented in strategy (ii), updates a fully connected classification head (linear probe) via iterative gradient methods during training and inference, which is subject to task-recency bias (Mai et al., 2021; Rymarczyk et al., 2023; Wang et al., 2023) and catastrophic forgetting (Zhang et al., 2023; Ramasesh et al., 2021). Class-prototype methods (Zhou et al., 2023b; Janson et al., 2022; McDonnell et al., 2023) that follow strategy (iii), in contrast, offer a promising alternative, as they utilize extracted features directly, making them more resource-efficient than full body adaptation and demonstrating greater robustness compared to PETL methods applied to the trainable linear probe.
@@
In this work, we hypothesize that representations are most effectively leveraged from multiple layers to construct a classifier based on first-order (class prototypes) and second-order (Gram matrix) feature statistics. Such a strategy is inspired by works on Neural Style Transfer (Gatys et al., 2016; Jing et al., 2020), which involve applying the artistic style of one image to the content of another image, a problem that can be viewed through the lens of domain adaptation (Li et al., 2017). To achieve neural style transfer, a model has to disentangle image information into content and style (e.g., textures, patterns, colors). While content information is expressed in terms of activations of features at different layers (which translates to calculating multi-layer class prototypes in our case), style information is expressed as correlations between features of different layers (which corresponds to calculating multi-layer Gram matrices).
@@
We thus propose Multi-Layer Universal Prototypes (LayUP), a class-prototype method for CL that is based on the aforementioned strategy to disentangle style and content information of images at multiple layers of a pre-trained network. An overview of our method is given in Fig. 3. LayUP obtains rich and fine-grained information about images to perform a more informed classification that is less sensitive to domain shifts. We additionally experiment with different parameter-efficient fine-tuning strategies to further refine the intermediate representations obtained by LayUP. Across various CL benchmarks and learning settings, our method enhances performance and narrows the gap to the upper bound by as much as 80% (Stanford Cars-196 in the CIL setting, cf. Sec. 5.2) compared to the next best baseline, while also reducing its memory and compute requirements by 81% and up to 90% (cf. Sec. 4.3), respectively. Beyond that, our method can serve as a versatile plug-in to enhance existing class-prototype methods, consistently yielding an absolute performance gain of up to 31.1% (cf. Sec. 5.7).
@@
Our contributions are threefold: (1) We present and examine in detail the CL strategy with pre-trained models and show why it benefits from extracting intermediate representations directly for classification. We further demonstrate the advantages of leveraging cross-correlations of features within and between multiple layers to decorrelate class prototypes. (2) Building on our insights, we propose LayUP, a novel class-prototype approach to CL that leverages second-order statistics of features from multiple layers of a pre-trained model. Inspired by prior works (Zhou et al., 2023a; McDonnell et al., 2023), we experiment with different methods for parameter-efficient model tuning and extend our method towards first session adaptation. Our final approach is not only conceptually simple, but also benefits from low memory and computational requirements, and integrates seamlessly with any pre-trained model. (3) We report performance improvements with a pre-trained ViT-B/16 (Dosovitskiy et al., 2021) backbone on the majority of benchmarks the Class-Incremental Learning (CIL) and Domain-Incremental Learning (DIL) settings (Van De Ven et al., 2022) as well as in the challenging Online Continual Learning (OCL) setting. We show that LayUP is especially effective under large distributional shifts and in the low-data regime. Our results highlight the importance of taking a deeper look at the intermediate layers of pre-trained models to better leverage their powerful representations, thus identifying promising directions for CL in the era of large foundation models. The source code is available at https://github.com/ky-ah/LayUP.
@@ 2.1 Continual learning
Traditional approaches to CL consider training a model from scratch on a sequence of tasks, while striking a balance between adapting to the currently seen task and maintaining high performance on previous tasks. They can be broadly categorized into the paradigms of regularization (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2018; Aljundi et al., 2018; Dhar et al., 2019), which selectively restricts parameter updates, (pseudo-)rehearsal (Rebuffi et al., 2017b; Chaudhry et al., 2019; Buzzega et al., 2020; Prabhu et al., 2020), which recovers data either from a memory buffer or from a generative model, and dynamic architectures (Rusu et al., 2016; Yoon et al., 2018; Serra et al., 2018), which allocate fully or partially disjoint parameter spaces for each task.
@@ 2.2 Continual learning with pre-trained models
As an alternative to the methods above that adjust the feature space via cross-entropy loss and backpropagation, constructing prototypes directly from the pre-trained features emerges as a cost-efficient alternative to overcome forgetting during CL. Janson et al. (2022) and Zhou et al. (2023b) find that a conceptually simple Nearest Mean Classifier (NMC) outperforms various prompt learning methods. ADAM (Zhou et al., 2023b) combines NMC with FSA and concatenates the features from the initial pre-trained ViT and the adapted parameters. Other methods leverage second-order feature statistics, i.e., covariance information, for class prototype accumulation: FSA+LDA (Panos et al., 2023) applies an incremental version of linear discriminant analysis to a pre-trained ResNet encoder (He et al., 2016). RanPAC (McDonnell et al., 2023) uses high-dimensional random feature projections to decorrelate class prototypes of a pre-trained ViT backbone.
All the aforementioned approaches have in common that they consider only the last layer representations (or, features) for classification. We provide a different perspective to the prototype-based approach to CL and show that intermediate representations can add valuable information to the linear transformation of class prototypes (cf. Eq. 6), which leads to better class separability and increased robustness to large distributional shifts.
@@ 3.1 Continual learning problem formulation
We consider a feature extractor $\phi(\cdot)$ composed of $L$ consecutive layers and a sequence of training datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$, where the $t$th task $\mathcal{D}_{t}=\{({\bm{x}}_{t,n},y_{t,n})\}_{n=1}^{N_{t}}$ contains pairs of input samples ${\bm{x}}_{t,n}\in\mathcal{X}_{t}$ and their ground-truth labels $y_{t,n}\in\mathcal{Y}_{t}$. The total number of classes observed in $\mathcal{D}$ (i.e., the number of unique elements in $\mathcal{Y}$) is denoted with $C$. Given an arbitrary input sample ${\bm{x}}\in\mathcal{X}$, the $d_{l}$-dimensional encoded features from the $l$th layer of model $\phi(\cdot)$ are represented as $\phi_{l}({\bm{x}})\in\mathbb{R}^{d_{l}}$. Thus, the features from the last (or representation) layer of the model are given by $\phi_{L}({\bm{x}})\in\mathbb{R}^{d_{L}}$.
Our focus in this work is rehearsal-free CL, where historical data cannot be fetched for replay. In the Class-Incremental Learning (CIL) setting, label spaces are partially disjoint between tasks. Conversely, in the Domain-Incremental Learning (DIL) setting, label spaces are kept across tasks, but samples within a label space are subject to distributional shifts. Online Continual Learning (OCL) is a more stringent definition of the aforementioned settings, where each datum in the stream of tasks is observed only once. All three CL settings prohibit the model from accessing task identifiers during testing; notably, our method does not require access to task-specific information during training either.
@@ 3.2 Class-prototype methods for CL with pre-trained models
To leverage the powerful representations from large-scale pre-training while overcoming the limitations of full body adaptation and linear probing (cf. Sec. 1), class-prototype methods extract and accumulate features from the last layer of a pre-trained model to construct representatives for each class. The most straightforward class-prototype method is the Nearest Mean Classifier (NMC), which aggregates for each class $y\in\mathcal{Y}$ training sample features via averaging to obtain class prototype $\overline{{\bm{c}}}_{y}$:
$\overline{{\bm{c}}}_{y}=\frac{1}{K}\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\mathbbm{1}(y=y_{t,n})\phi_{L}({\bm{x}}_{t,n}),$ (1)
with $\mathbbm{1}(\cdot)$ denoting the indicator function and $K=\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\mathbbm{1}(y=y_{t,n})$.
During inference, NMC selects for each test sample’s feature vector the class with the smallest Euclidean distance (Janson et al., 2022) or highest cosine similarity (Zhou et al., 2023b) with its class prototype. Considering the cosine similarity measure and some ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, the predicted class label is obtained by:
@@ 4 LayUP
Figure 4: Classification performance of intermediate layers. Comparison for two different pre-training paradigms of a ViT-B/16 (Dosovitskiy et al., 2021) backbone and split ImageNet-R (left) and CIFAR-100 (right) datasets. For each intermediate layer $l\in\{1,\dots,L-1\}$ (where $L=12$ denotes the final representation layer), the bars represent the percentage of classes for which a classifier, utilizing Eq. 3 at layer $l$, surpasses the accuracy of the classifier at the $L$th layer.
@@ 4.1 Class prototyping from second-order intermediate features
We argue that a combination of (i) enriching the last layer representations with hierarchical multi-layer features and (ii) decorrelating multi-layer features—which represent image properties such as content and style—via Gram matrix transformation increases robustness to domain shifts and thus improves generalizability to downstream continual tasks. To this end, we propose to integrate embeddings from multiple layers via concatenation into a ridge regression estimator as defined in Eq. 3. Let
@@
denote the concatenated input features of some input sample ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, extracted from the last $k$ layers of the pre-trained model $\phi(\cdot)$. Such features can be, e.g., classification token embeddings ([CLS]) of a transformer (Vaswani et al., 2017) encoder or flattened feature maps of a ResNet (He et al., 2016) encoder. This yields a modified estimator
@@
with $d_{(k)}=(d_{L-k+1}+\dots+d_{L-1}+d_{L})$-dimensional identity matrix ${\bm{I}}$ and
@@
We first motivate our approach by showing that intermediate representations capture expressive class statistics for prototype-based classification. For the split CIFAR-100 (Krizhevsky, 2009) and ImageNet-R (Hendrycks et al., 2021a) datasets and two ViT models, we construct one prototype-based classifier using Eq. 3 per layer $l\in\{1,\dots,L-1\}$ and measure the percentage of classes that layer $l$ predicts better than the last layer $L$ ($L=12$ for ViT-B/16). The results are provided in Fig. 4. Across both datasets, the classifiers from the last five intermediate layers outperform the representation layer classifier in up to 32% of the classes. The advantage of multi-layer classifiers is more pronounced in the ViT model pre-trained on 21 000 different ImageNet classes and additionally fine-tuned on 1 000 different ImageNet classes (IN1K) compared to the ViT model only pre-trained on 21 000 different ImageNet classes (IN21K). The results indicate that representations derived from intermediate layers generally provide meaningful knowledge to more effectively distinguish classes at various hierarchical levels.
@@
Figure 5: Comparison of techniques to integrate intermediate representations. LayUP implementations for different values of $k$ using a shared representation and Gram matrix as in Eq. 6 versus averaging over separate ridge (or, Gram) classifiers for each layer using Eq. 3. Results are reported as average accuracy scores over CL training on split ImageNet-R (left) and CIFAR-100 (right) datasets following phase B of Alg. 1.
@@
There are two intuitive ways to integrate intermediate representations into the class-prototype classifier: Averaging over $k$ separate classifiers per Eq. 3 or concatenating representations from the last $k$ layers to obtain shared class prototypes per Eq. 6. Fig. 5 shows average accuracies for two split datasets and different values of $k$. For each $k$, transforming shared class-specific information via Gram matrix inversion, as described in Eq. 6, proves to be superior to averaging across separate classifiers. This finding indicates that the correlations across layers, captured in the shared Gram matrix (cf. Eq. 7), add meaningful information to enhance class separability.
@@
While performance generally benefits from higher values of $k$, the difference in performance between $k=4$ and $k=6$ is more pronounced than that between $k=6$ and $k=12$. Considering that a higher $k$ corresponds to increased computational and memory demands as $\mathbf{G}$ and $\mathbf{c}_{y}$ increase in dimension, this substantiates $k=6$ as a reasonable choice for maximum representation depth. Although concatenating representations from multiple consecutive layers results in higher memory demands compared with using a final-layer classifier or averaging over layer-wise classifiers, we will demonstrate in Sec. 4.3 that our method remains significantly lighter in memory and computational requirements than other competitive class-prototype methods for CL.
@@
Interestingly, the performance of the final-layer classifier $k=1$, as illustrated in Fig. 5, is approximately equivalent to that achieved by averaging over separate classifiers for $k=4$ and $k=6$. This observation challenges the argument that class embeddings from multiple consecutive layers, which are individually strong at predicting specific classes (cf. Fig. 4), might introduce noise that impairs performance when combined via averaging.